what is the kinetic energy in J of a 60.0 g tennis ball approaching a tennis racket at 20.0 m/s
Answer #1
mv^2/2
=0.06*20^2/2
=12 J
Answer #2
12 j
Answer #3
As we know

KE =1/2mv2

=1/2*60*400

=12000J